How many units of horizontal shifting occurs?
from: f(x) = 2tan(x/3 + 1) + 2
to: f(x) = 2tan(x/3 + 4) + 2

Answer by Nate(3500)

Generally, you would subtract the two indicators for horizontal placement.
f(x) = a*tan(bx + c) + d
pi(c(of the first) - c(of the second)) = horizontal displacement
pi(1 - 4) = -3pi or about -9.42477796....
The graph is moved about nine and a half units to the left.